OpenStudy (anonymous):

Someone with vector knowledge! Please! Find the vector and parametric equations for this line: goes through (7,5) with inclination 45º.

OpenStudy (anonymous):

Ok, we'll give it a go.... So our line (in ordinary slope-intercept form) is y=x-2. We together on this so far?

OpenStudy (anonymous):

Parameterize x=t, y=t-2. Not very imaginative, but I think it works for any value of x,y, and t.

OpenStudy (anonymous):

If I want to put it into a vector, you could look at it like\[V(vector)=t *i(unit vector)+(t-2)*j(unitvector)\]I'm sorry, my mastery of LATEX is just about non-existent.

OpenStudy (anonymous):

The V(vector) is a function of parameter t.

OpenStudy (anonymous):

Hope I was helpful.

OpenStudy (mertsj):

x=7-7t y=5-7t Vector Equation: (7,5)+(-7,-7)t